OpenStudy (anonymous):

Need some help putting these critical values into interval form: x = 0, -sqrt(6/5), sqrt(6/5)

OpenStudy (anonymous):

should be something like this, correct? \[(-\infty, -\sqrt{6/5}) U (-\sqrt{6/5}, 0) U(0, \sqrt{6/5}) U(\sqrt{6/5}, \infty)\]